what is the measure of the larger angle of two supplementary angles if two-fifths of the larger angle is 5 less than the smaller angle?
Answer by nerdybill(7384)

Supplementary: two angles whose sum is 180
.
Let x = larger angle
180-x = smaller angle
.
(2/5)x = (180-x)-5
(2/5)x = 175-x
2x = 875 - 5x
7x = 875
x = 125 degrees